OpenStudy (cookiemonster18):

1. 3(x+4)=-2(x+5) 2.5x+9-7x-10=3x+4(x+2)

OpenStudy (harvey03):

1. = x + 2

OpenStudy (cookiemonster18):

whats X?? for #1 I don't understand how to solve these problems

OpenStudy (harvey03):

It's called expanding brackets. You do 3 times x which = 3x, then 3 times 4 which = 12. Then you do -2 times x which = -2x, and then -2 times 5 which = -10. Finally you add all of these factors together, giving you an answer of x + 2

OpenStudy (taemin):

\[3x+12=-2x-10\] try this

OpenStudy (cookiemonster18):

@harvey03 can you help with #2

OpenStudy (harvey03):

I am not too sure how to do that one, sorry

OpenStudy (taemin):

\[5x+9-7x-10=3x+4(x+2) do like terms \]

OpenStudy (taemin):

\[5x+9-7x-10 do this half first\]

OpenStudy (cookiemonster18):

how if I don't know what x is??? @taemin

OpenStudy (taemin):

hold on i'm going to break it down

OpenStudy (cookiemonster18):

okay @taemin

OpenStudy (cookiemonster18):

you there @taemin

OpenStudy (taemin):

yep

OpenStudy (taemin):